A store marks up items to make 89% profit. If an item costs $3.50 from the supplier, what will the price be?
Answer by Fombitz(32388)

Store Cost + Profit = Customer Cost
Store Cost + 0.89*(Store Cost) = Customer Cost
1.89*(Store Cost) = Customer Cost
Customer Cost = 1.89*($3.50)
Customer Cost = $6.62 (rounded up from 6.615).